# Topology of dyonic AdS black holes with quasitopological electromagnetism in Einstein-Gauss-Bonnet gravity

### v2

@@ IV Nature of Critical points
Down panel: We can find that the isobaric curves $(P_{CP3}$=$0.03308(purple)$; $P$=$0.003309(orange)$; $P_{CP2}$=$0.003311(blue)$; $P=0.025(green);(P_{CP1}=0.050611(red))$ exhibit temperature extreme points. Similarly, the green dots ($CP_{1}$ and $CP_{2}$) correspond to the point of maximum extreme temperature point, while the blue dot (novel critical point $CP_{3}$) accurately corresponds to the point of minimum extreme temperature point. Based on research findings cq3 , it has been observed that there is no occurrence of a first-order phase transition in close proximity to the conventional critical point $CP_{2}$, which can also be found in we2 ; we7 . In FIG.10, we give the temperature $(T_{p}=0.05328)$ at which the $P=0.025$ black hole undergoes a phase transition. Notably, the application of Maxwell’s law of equal area can effectively remove the unstable region near $CP_{2}$. However, it is not suitable for indicating a first-order phase transition. Therefore, an alternative classification should be employed, it can be inferred that $CP_{2}$ may serve as a phase annihilation point.
V Conclusions
In summary, we employ the Duan ’s $\varphi$-mapping theory to examine the topological structure of thermodynamics in black holes. The categorization of critical points for black holes is carried out across diverse dimensions of space-time. In the scenario where $d=6$, the black hole will have triple points when $\alpha$ specific value is assumed by the GB coupling constant. Our focus lies on assessing how this coupling constant impacts the critical point. This investigation contributes to a more profound comprehension of EGB modified black holes. The primary highlights can be summarized as follows:
$(i)$ Black holes have a critical point in any space-time dimension, the presence of a small/large BH phase transition gives rise to a conventional (physical) critical point, attributed to the total topological charge $-1$. Different $d$ will lead to different positions of the critical point in vector field $n$. Initially, as d increases, the critical point also increases before eventually decreasing again in TABLE I.
@@ V Conclusions
$(ii)$ For $d=6$, we will find that when the coupling constant satisfies $\alpha=0.5$ and $1$, the system will appear triple points, but the phase structure is completely different. More specifically, the three critical points can be classified into two conventional (physical) critical points, $(CP_{1}$ and $CP_{2})$, as well as a novel critical point $(CP_{3})$, the novel critical point is not a physical point because it does not minimize the Gibbs free energy. When a = 1, the system has only one traditional ( physical ) critical point $(CP_{1})$. In both cases, the total topological charge is $-1$.
@@
$(iii)$ From the isobaric curve and critical temperature in FIG.7, we can observe that the conventional critical point is located at the maximum extreme point of temperature, and the unstable region in close proximity to it can be eliminated through the application of Maxwell’s equal area law. It is found that the critical point $CP_{3}$ is no longer applicable. Noted that the novel critical point appears at the minimum extreme point of temperature. The results also reveal a decrease in the number of phases along the isobaric curve at the conventional critical point as pressure increases $(\alpha=1)$, while the number of phases increases at the novel critical point. Therefore, we define the $CP_{2}$ as the phase annihilation point $(\alpha=1)$.
@@ Acknowledgements.
This work is supported by the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2022]-314), and the National Natural Science Foundation of China (Grant Nos. 12265007 and 12264061).Bekir Can Lütfüoğlu is grateful to Excellence project PřF UHK 2211/2023-2024 for the financial support.

### v3

@@ IV Nature of Critical points
Similarly, the green dots ($CP_{1}$ and $CP_{2}$) correspond to the point of maximum extreme temperature point, while the blue dot (novel critical point $CP_{3}$) accurately corresponds to the point of minimum extreme temperature point. Based on research findings cq3 , it has been observed that there is no occurrence of a first-order phase transition in close proximity to the conventional critical point $CP_{2}$, which can also be found in o1 . In FIG.8, we give the temperature $(T_{p}=0.05328)$ at which the $P=0.025$ black hole undergoes a phase transition. Notably, the application of Maxwell’s law of equal area can effectively remove the unstable region near $CP_{2}$. However, it is not suitable for indicating a first-order phase transition. Therefore, an alternative classification should be employed, it can be inferred that $CP_{2}$ may serve as a phase annihilation point.
V Dyonic AdS black holes as a topological defect
In this section, we intend to examine the topological number and defect curves pertaining to the dyonic black hole featuring QTE within EGB gravity. By incorporating the equations for mass and entropy (Eq. 12 and Eq. 16) into Eq. 31, we can establish the generalized free energy
$\mathcal{F}=\frac{1}{32\left(3-4d+d^{2}\right)\pi}r_{+}^{-5-d}\mathcal{V}_{d-2}$ (40)
×$\quad\times\left[(d-1)Q_{m}^{2}r_{+}^{8}+2(d-3)r_{+}^{2d}\right.$
×$\quad\times\left(\left(2-3d+d^{2}\right)r_{+}^{2}\right.$
$\left.\quad+\left(24-50d+35d^{2}-10d^{3}+d^{4}\right)\alpha+16\pi r_{+}^{4}P\right)$
$\quad+(d-1)Q_{e}^{2}r_{+}^{8}{}_{2}F_{1}$
×$\left.\quad\times\left(1,\frac{d-3}{2(d-2)},\frac{7-3d}{4-24},-8\beta Q_{m}^{2}r_{+}^{4-2d}\Gamma(d-1)^{2}\right)\right]$
$-\frac{r_{+}^{d-2}\pi^{\left(\frac{d-1}{2}\right)}\left(1+2(d-3)(d-2)\alpha r_{+}^{-2}\right)}{2\tau\Gamma\left(\frac{d-1}{2}\right)},$
the components of the vector $\phi$ is given by
$\phi^{r_{+}}$ $=\frac{1}{16\Gamma\left(\frac{d-1}{2}\right)}\pi^{\frac{(d-3)}{2}}r_{+}^{-6-d}\left(-r_{+}^{8}Q_{m}^{2}+r_{+}^{2d}\right.$ (41)
×$\times\left(32P\pi r_{+}^{4}+2(-3+d)(-2+d)\right.$
×$\times\left(r_{+}^{2}+(d-5)(d-4)\alpha\right)$
$-\frac{8(d-2)\pi r_{+}\left(r_{+}^{2}+2(d-4)(d-3)\alpha\right)}{\tau}$
$\left.\left.-\frac{r_{+}^{8}Q_{e}^{2}}{r_{+}^{2d}+8r_{+}^{4}\beta\Gamma(d-2)^{2}Q_{m}^{2}}\right)\right),$
$\phi^{\Theta}=-\cot\Theta\csc\Theta.$ (42)
The expression for $\tau$ can be obtained by setting $\phi^{r}+$ equal to zero
$\tau=\frac{8(d-2)\pi r_{+}^{2d+1}\left(r_{+}^{2}+2\tilde{\alpha}\right)\chi_{2}}{-r_{+}^{8}Q_{m}^{2}\chi_{2}+r_{+}^{2d}\left(-r_{+}^{8}Q_{e}^{2}+2\left(\chi_{3}r_{+}^{2}+(-5+d)\chi_{1}\right)\chi_{2}\right)},$ (43)
here
$\chi_{2}=r_{+}^{2d}+8r_{+}^{4}\beta\Gamma(d-2)^{2}Q_{m}^{2},$ (44)
and
$\chi_{3}=6+(-5+d)d+16P\pi r_{+}^{2}.$ (45)
Next, we will detect the topological number of black holes in different dimensions.
V.1 The small/large phase transition $(d=5,7,8,9)$.
Since black holes $(d=5,7,8,9)$ only have small / large phase transition, accompanied by a critical point, we only study the case of $d=5$. In this case, the pressure is below its critical threshold, and similar to the phase transitions observed in van der Waals liquid/gas systems, small/large phase transitions occur within the black hole. To ascertain the zero points, we consider the unit vectors by setting $\Theta=\pi/2$ in $n^{1}$ and equating it to zero. By considering $\tau/r_{0}=15$, the unit vector $n$ is depicted on the $\left(r_{+}/r_{0},\Theta\right)$ plane in FIG.9, we find the presence of a zero point at $r_{+}/r_{0}=34.0601$. The vector changes its direction once it passes through a zero point along $\Theta=\pi/2$.By examining the zero point $ZP_{1}$, we observe that the winding number of $ZP_{1}$, the sum of these winding numbers yields a topological number $W=+1$.If we consider $\tau/r_{0}=20$ in FIG.10, we find that three zero points are located in $r_{+}/r_{0}=1.30526$,2.5324, 34.0601.We observe that the winding number of $ZP_{2}$ and $ZP_{4}$ is positive $(+1)$ while that of $ZP_{3}$ is negative $(-1)$. The sum of these winding numbers yields a topological number $W=+1-1+1=+1$. In FIG. 11, we present a single zero point vector diagram when $\tau/r_{0}=38$. Similarly, the same results appear when $d=7,8,9.$, the system shares a similar topological classification as the charged RN-AdS black holes we1 . The relationship between $r_{+}$ and $\tau$ is presented in FIG.12, as previously derived in Eq.36, revealing the existence of two turning points. In we1 , the authors established an annihilation point and a generation point by considering the condition $\frac{d^{2}\tau}{dr_{+}^{2}}>0$ and $\frac{d^{2}\tau}{dr_{+}^{2}}<0$ at a specific point $\tau_{c}$, respectively. Based on this, we can observe the presence of a point of generation and a point of annihilation.
Figure 9: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.001$. The black dot represents zero point ($(\tau/r_{0}=15,\alpha/r_{0}=0.5,\beta/r_{0}=0.1)$.
Figure 10: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.001$. The black dot represents zero point ($(\tau/r_{0}=20,\alpha/r_{0}=0.5,\beta/r_{0}=0.1)$.
Figure 11: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.01$. The black dot represents zero point ($(\tau/r_{0}=38,\alpha/r_{0}=0.5,\beta/r_{0}=0.1)$.
Figure 12: Defect curves in the $r_{h}/r_{0}-\tau/r_{0}$ plane with $Pr_{0}^{2}=0.001,\alpha/r_{0}=0.5,\beta/r_{0}=0.1$.
V.2 The triple points $(d=6)$.
Now, we will involve examining the topological number pertaining to the dyonic black hole featuring $\alpha/r_{0}=0.5$ and $\beta/r_{0}=0.1$. For the sake of simplicity, we consider three cases where the pressure $Pr^{2}_{0}$ is equal to 0.006, 0.0064, and 0.05. Here, $r_{0}$ represents an arbitrary length scale determined by a cavity that encompasses the black hole.
$(i)$ $Pr^{2}_{0}=0.006$
We find that this result is consistent with the results of other dimensional black holes $(d=5,7,8,9)$ in Figs.13-14. Under the background of $\Theta=\pi/2$, there are three zero points located at $r_{+}/r_{0}=1.08325,2.54691$, and $4.87928$, the sum of these winding numbers yields a topological number $W=+1-1+1=+1$. In addition, we can observe the presence of a point of generation and a point of annihilation.
$(ii)$ $Pr^{2}_{0}=0.0064$
Figure 13: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.006$. The black dot represents zero point ($(\tau/r_{0}=15.6)$.
Figure 14: Defect curves in the $r_{h}/r_{0}-\tau/r_{0}$ plane with $Pr_{0}^{2}=0.006$.
Figure 15: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.0064$. The black dot represents zero point ($(\tau/r_{0}=15.468)$.
Here, we adjust the pressure to achieve a value of $Pr_{0}^{2}=0.0064$, the additional features related to defects and defect curves will be observed. We observe that the $\left(r_{+}/r_{0},\Theta\right)$ plane in FIG.15 exhibits five zero points at $r_{+}/r_{0}=1.29839,1.51839,1.88713,2.40912$, and $3.83608$, as indicated by the unit vector $n$. Similarly, we can observe that the zero points $ZP_{4}$, $ZP_{6}$, and $ZP_{8}$ possess a winding number of $+1$, whereas $ZP_{5}$ and $ZP_{7}$ exhibit a winding number of $-1$. By summing up all the winding numbers, the topological number to be $W=+1-1+1-1+1=+1$.
Figure 16: Defect curves in the $r_{h}/r_{0}-\tau/r_{0}$ plane with $Pr_{0}^{2}=0.0064$.
Interestingly, it can be observed that the system exhibits two points of generation and two points of annihilation in FIG.16, this implies that with an increase in pressure, there will be alterations in both the generation point and the annihilation point. The total topological number, however, remains unchanged at $Pr_{0}^{2}=0.006$. For dyonic black holes with QTE, the $r_{h}/r_{0}-\tau/r_{0}$ plane manifests two separate curves, ni2 . However, our findings indicate that the continuity of the curve in the $r_{h}/r_{0}-\tau/r_{0}$ plane is evident, this implies that the characteristics of $r_{h}/r_{0}-\tau/r_{0}$ are influenced by the coupling constants $\alpha$ and $\beta$.
$(iii)$ $Pr^{2}_{0}=0.05$
For this case, the pressure exceeds its critical threshold, resulting in the absence of any phase transition. We depict the unit vector in FIG.17 when $\tau/r_{0}=2$. It is evident that the zero point exists at coordinates $\left(\frac{r_{h}}{r_{0}},\Theta\right)=(2.01201,\frac{\pi}{2})$. The winding number of the zero point $ZP_{9}$ is +1. According to FIG.18, it can be observed that the system does not produce any generation points or annihilation points.
Figure 17: Unit vector $n=\left(n^{1},n^{2}\right)$ shown in $\Theta$ vs $r_{+}/r_{0}$ plane for $\operatorname{Pr}_{0}^{2}=0.05$. The black dot represents zero point ($(\tau/r_{0}=10)$.
Figure 18: Defect curves in the $r_{h}/r_{0}-\tau/r_{0}$ plane with $Pr_{0}^{2}=0.05$.
Of specific significance, the defects observed on this curve exhibit a winding number of $+1$, resulting in a topological number $W=+1$. Therefore, the dyonic AdS black holes with QTE in EGB gravity belong to the same topological class as the charged RN-AdS black holes we1 .
VI Conclusions
In summary, when considering a particular value for the GB coupling constant $\alpha$, the phase structure of 6-dimensional dyonic AdS black holes becomes more intricate, with the appearance of triple points cq3 . We initially investigate the topological charge associated with the critical point, followed by an analysis of dyonic AdS black holes as a thermodynamic space’s topological defect, in order to assess the topological structure of dyonic AdS black holes. The primary highlights can be summarized as follows:
@@ VI Conclusions
$(i)$ We find that when the coupling constant satisfies $\alpha=0.5$ and $1$, the system will appear triple points, but the phase structure is completely different. More specifically, the three critical points can be classified into two conventional (physical) critical points, $(CP_{1}$ and $CP_{2})$, as well as a novel critical point $(CP_{3})$, the novel critical point is not a physical point because it does not minimize the Gibbs free energy. When a = 1, the system has only one traditional ( physical ) critical point $(CP_{1})$. In both cases, the total topological charge is $-1$.
@@
$(ii)$ From the isobaric curve and critical temperature in FIG.4, we can observe that the conventional critical point is located at the maximum extreme point of temperature, and the unstable region in close proximity to it can be eliminated through the application of Maxwell’s equal area law. It is found that the critical point $CP_{3}$ is no longer applicable. Noted that the novel critical point appears at the minimum extreme point of temperature. The results also reveal a decrease in the number of phases along the isobaric curve at the conventional critical point as pressure increases $(\alpha=1)$, while the number of phases increases at the novel critical point. Therefore, we define the $CP_{2}$ as the phase annihilation point $(\alpha=1)$.
$(iii)$ We regard dyonic AdS black holes as a topological defect in the thermodynamic space, wherein the black hole has a generation point and annihilation point upon the pressure being below critical threshold. The black hole exhibits two points of generation and annihilation as the pressure increases. When the pressure surpasses a critical threshold, there are no production or annihilation points. However, the total topological number of black holes in different dimensions is $1$, the system shares a similar topological classification as the charged RN-AdS black holes we1 .
@@ Acknowledgements.
We would like to thank the anonymous referees for their valuable comments on improving our paper. This work is supported by the Doctoral Foundation of Zunyi Normal University of China (BS [2022] 07, QJJ-[2024]-203), and the National Natural Science Foundation of China (Grant Nos. 12265007). Also, H. H. is grateful to Excellence project FoS UHK 2203/2025-2026 for the financial support. B.C.L is grateful to Excellence project FoS UHK 2212 / 2021-2022 for the financial support.